OpenStudy (anonymous):

Need help with Calc1 - Integrals

OpenStudy (anonymous):

F(x)=\[\int\limits_{\cos x}^{5x} \cos(u^{2})du\]

OpenStudy (anonymous):

Find derivative I got cos(5x)^2 *5+-cos(cosx)^2*-sinx

OpenStudy (anonymous):

You need to use the fundamental theorem of calculus here: \[F'(x)=5\cos(25x^2)-\cos(\cos^2x)(-\sin x)=5\cos(25x^2)+\sin(x)\cos(\cos^2x).\]

OpenStudy (anonymous):

You can derive a general formula given by: \[{d \over dx}\int\limits_{g(x)}^{h(x)}f(x)dx=f(h(x))h'(x)-f(g(x))g'(x).\]

OpenStudy (anonymous):

our answers are the same

OpenStudy (anonymous):

That's good.

OpenStudy (anonymous):

:) thanks then I'll ask another maybe u'll get to use ur brain a little more

OpenStudy (anonymous):

Lol! Ask away!!

OpenStudy (anonymous):

Evaluate the Integral : \[\int\limits_{-1}^{2}x-2|x|dx\]

OpenStudy (anonymous):

No idea what to do with the |x|

OpenStudy (anonymous):

I'm sure you know how to do the first term, so I'll focus on the integration of |x| from -1 to 2. To integrate the absolute value of x you need to redefine the function. You know that \(|x|=-x \text{ for } x<0\), so we can write the integral as: \[\int\limits_{-1}^{2}|x|dx=\int\limits_{-1}^{0}-xdx+\int\limits_0^2xdx\]

OpenStudy (anonymous):

true

OpenStudy (anonymous):

Evaluate it and then multiply it by the constant \(-2\) and add to it the value you get from integrating the first term.

OpenStudy (anonymous):

Makes sense?

OpenStudy (anonymous):

lemme try

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

\[(1/2)x^{2}-3\] then i plug in each and do F(b)-F(a) right?

OpenStudy (anonymous):

I didn't solve it actually, but you can always check your answers using wolfarmalpha :D

OpenStudy (anonymous):

I don't know where the -3 came from though.

OpenStudy (anonymous):

integrate x ---> (1/2)x^2 then i pluged in the numbers and added them together

OpenStudy (anonymous):

If you plugged in the numbers, why do I still see x in there?

OpenStudy (anonymous):

thats the first term the -3 is the -2|x| part

OpenStudy (anonymous):

Oh I see. You're right then, you just need to plug -1 to 2 in 1/2 x^2.

OpenStudy (anonymous):

Are you sure about the -3? It should be -5.

OpenStudy (anonymous):

lemme check ur prob right

OpenStudy (anonymous):

ehhh im getting (3/2) * -2 = -6/2 = -3

OpenStudy (anonymous):

\[\int\limits_{-1}^{2}\left| x \right|dx=\int\limits_{-1}^0-xdx+\int\limits_0^2xdx={-x^2 \over 2}|_{-1}^0+{x^2 \over 2}|_0^2=(0-{-1 \over 2})+({4 \over 2}-0)={5 \over 2}\]

OpenStudy (anonymous):

\(\large =\frac{5}{2}\)

OpenStudy (anonymous):

my bad forgot to do the 0 minus on the negative one!

OpenStudy (anonymous):

I thought so!

OpenStudy (anonymous):

My appologies so now its easy \[x^{2}/2-5 \] [-1, 2] ---> 4/2=2-5=-3 ( 1/2)-5 = -9/2 so -3-(-9/2) = 3/2

OpenStudy (anonymous):

Wait! Not like this. Do first \(\frac{x^2}{2}|_{-1}^2=\frac{4}{2}-\frac{1}{2}=\frac{3}{2}\). So the final value of our integral is \(\frac{3}{2}-5=-\frac{7}{2}.\)

OpenStudy (anonymous):

I gotta go now!

OpenStudy (anonymous):

sorry for being difficult thanks!

OpenStudy (anonymous):

no problem! glad to help :D